You own a stock that had returns of 10.70 percent, −15.50 percent, 19.80 percent, and 18.60 percent over the past four years. What was the arithmetic average return for this stock?

Multiple Choice

  • 9.10%

  • 8.74%

  • 8.40%

  • 7.37%

  • 7.89%
Answer #1

Arithmetic average return of stock

=( X1+X2+X3+.....+XN)/N

= (10.70 - 15.50 +19.80+18.60)/4 = 8.40%